# A Bayesian Generative Modeling Approach for Arbitrary Conditional Inference

### v1

@@ Abstract
Modern data analysis increasingly requires flexible conditional inference $P(X_{\mathcal{B}}|X_{\mathcal{A}})$ where $(X_{\mathcal{A}},X_{\mathcal{B}})$ is an arbitrary partition of observed variable $X$. Existing conditional inference methods lack this flexibility as they are tied to a fixed conditioning structure and cannot perform new conditional inference once trained. To solve this, we propose a Bayesian generative modeling (BGM) approach for arbitrary conditional inference without retraining. BGM learns a generative model of $X$ through an iterative Bayesian updating algorithm where model parameters and latent variables are updated until convergence. Once trained, any conditional distribution can be obtained without retraining. Empirically, BGM achieves superior prediction performance with well calibrated predictive intervals, demonstrating that a single learned model can serve as a universal engine for conditional prediction with uncertainty quantification. We provide theoretical guarantees for the convergence of the stochastic iterative algorithm, statistical consistency and conditional-risk bounds. The proposed BGM framework leverages the power of AI to capture complex relationships among variables while adhering to Bayesian principles, emerging as a promising framework for advancing various applications in modern data science. The code for BGM is freely available at https://github.com/liuq-lab/bayesgm.
@@ 1 Introduction
Conditional inference concerns the problem of inferring the distribution of a response variable given a set of predictor variables, which is a fundamental task in both machine learning and statistical science (Reid, 1995). Numerous machine learning methods have been developed for estimating the mean response conditioning on the predictor variables, which is widely known as the regression problem. However, the conditional mean alone may not always be sufficient in many scenarios where different distributional characteristics or even the full conditional distribution are necessary. For example, accessing to the full conditional distribution enables the construction of prediction intervals and the quantification of predictive uncertainty, which are critical in high-stakes domains such as finance, medicine, and risk management, where decisions must account for variability rather than rely solely on point estimates (Sullivan, 2015; Soize, 2017).
A wide range of statistical and machine learning methods have been developed for estimating conditional distributions. Classical approaches include kernel and local-likelihood estimators, series and spline expansions, and mixture or tree-based models, which approximate conditional densities in a fully nonparametric manner (Fan et al., 1996; Hyndman et al., 1996; Stone, 1994; Bishop, 1994; Meinshausen and Ridgeway, 2006). While these methods provide valuable theoretical insight and flexibility, they suffer significantly from the curse of dimensionality where the sample size required for accurate estimation grows exponentially with the number of conditioning variables (Stone, 1980). As a result, their empirical performance deteriorates rapidly in high-dimensional settings. In addition to statistical inefficiency, many nonparametric estimators are computationally intensive, making them difficult to be widely applied to modern large-scale data where both the sample size and the feature dimension are substantial.
Recent years have witnessed rapid progress in modern conditional density estimation, driven by advances in generative artificial intelligence (AI). Neural conditional density estimators extend classical ideas by parameterizing conditional distributions through neural networks, as in mixture density networks, autoregressive models, and conditional normalizing flows (Errica et al., 2021; Papamakarios et al., 2017; Kingma and Dhariwal, 2018; Melnychuk et al., 2023). These approaches can capture highly nonlinear relationships and complex, multimodal conditionals that are intractable for traditional nonparametric estimators. Despite their expressive power, these models are typically designed for a fixed conditioning structure. For example, predicting a particular subset of variables given another and often require retraining or reconfiguration when the conditioning set changes. Moreover, they focus primarily on accurate density estimation or conditional generation, with limited mechanisms for calibrated uncertainty quantification.
@@
Parallel to advances in conditional density estimation, a separate line of research has been focused on conformal prediction (CP), which provides a model-agnostic framework for constructing prediction sets with guaranteed finite-sample coverage (Vovk et al., 2005; Lei et al., 2018; Barber et al., 2021). Conformal prediction methods transform point predictors into valid predictive intervals or regions without relying on any parametric assumption, ensuring that the true response lies within the constructed set with a user-specified probability. Recent developments have substantially expanded the scope of conformal prediction methods, including quantile-based (Romano et al., 2019), as well as localized conformal prediction (Guan, 2023) that adapts coverage to heterogeneous data distributions. These CP approaches have become a cornerstone for uncertainty quantification in modern machine learning applications due to their simplicity and strong theoretical guarantees (Kato et al., 2023). However, conformal prediction typically offers marginal coverage rather than full conditional calibration, and the resulting prediction sets depend on a specific base model trained and are also constrained to a fixed conditioning structure.
@@
To address these challenges, we propose a Bayesian Generative Modeling (BGM) framework that leverages the representational power of modern AI techniques while adhering to the foundational principles of Bayesian inference. BGM learns the underlying generative process of the observed variables from a low-dimensional latent space, estimated through a stochastic iterative updating algorithm. Once trained, BGM enables arbitrary conditional inference, predicting the distribution of any subset of variables given the remaining without retraining or modifying the model architecture. This “train once, infer anywhere” property allows BGM to naturally accommodate dynamically changing conditioning sets that commonly arise in practice, such as varying missing-data patterns, multimodal prediction tasks, and cross-domain imputations. By maintaining a coherent Bayesian formulation, BGM delivers calibrated predictive intervals and principled uncertainty quantification at any user-specified significance level. The proposed BGM framework thus combines the flexibility of modern AI models with the statistical coherence of Bayesian inference, offering a powerful and scalable solution for conditional prediction and uncertainty quantification in high-dimensional settings. We summarize the contributions of BGM as follows.
  • To the best of our knowledge, BGM is the first statistical framework for arbitrary conditional inference across any partition of the variables with a single trained model, providing both point estimates and predictive intervals. This capability fundamentally extends existing conditional inference methods, which are typically limited to one fixed conditioning structure or require separate models for different conditional queries.
  • BGM provides significant flexibility and scalability since the iterative algorithm only requires a mini-batch of data at each step to learn the data generative process from a low-dimensional latent space. This design allows BGM to efficiently handle large, high-dimensional datasets while retaining the theoretical coherence of Bayesian inference. The framework naturally accommodates modern deep network architectures, enabling practitioners to adapt BGM to a wide range of tasks.
@@
  • Extensive experiments demonstrate that BGM achieves state-of-the-art predictive accuracy and calibrated uncertainty quantification, consistently outperforming leading conformal prediction methods. These results highlight BGM’s practical value as a versatile and principled tool for conditional prediction and uncertainty assessment in modern data science applications.
@@ 2.1 Problem Setup
Although estimating the expected response $\mathbb{E}[X_{\mathcal{B}}|X_{\mathcal{A}}]$ remains central to most regression analysis, this statistic alone provides an incomplete picture of the underlying uncertainty. Practitioners often require much richer distributional information, such as conditional variances, quantiles, tail probabilities, or predictive intervals, to assess risk, reliability, and variability in predictions. Therefore, our focus extends beyond point estimation to the broader goal of accurately characterizing the full conditional distribution of interest, providing both point estimates and uncertainty qualifications across arbitrary subsets of variables to condition on.
@@ 2.2 Generative Process of Observed Data
where $\theta$ is a common parameter to specify all the conditional distribution of $X$ given $Z$. The prior distributions of both $Z$ and $\theta$ are set to be multivariate normal distributions. By default, we model the conditional distribution as normal distributions for continuous variables and logistic regression for discrete variables. In a typical setting, the generative processes are defined as follows
@@
where both mean and covariance matrix are learnable functions of latent variable $Z$ parameterized by $\theta$. In practice, we can further simplify the covariance matrix $\Sigma(Z)$ as a digonal structure $\Sigma(Z)=diag(\sigma^{2}_{1}(Z),...,\sigma^{2}_{p}(Z))$. In this diagonal simplification, $X_{\mathcal{A}}$ is independent of $X_{\mathcal{B}}$ given the latent variable $Z$. A richer covariance structures can capture residual conditional dependence (see our Discussion). Note that the generative function $G$ provides both mean function $\mu(Z)$ and covariance function $\Sigma(Z)$, which is represented by a neural network parametrized by $\theta$. To account for uncertainty or variation in model parameter space, we can also adopt Bayesian neural network (BNN), where model parameters are treated as random variables in stead of deterministic values (Goan and Fookes, 2020; Jospin et al., 2022).
@@ 2.3 Stochastic Iterative Updating Algorithm
where $C_{1}=C-\frac{p}{2}log(2\pi)$, $X_{i}$ is the $i$-th element in $X$, and $\mu_{i}(Z)$ is the $i$-th element in the mean vector $\mu(Z)$. We update latent variable $Z$ for each sample during the training stage by maximizing (4) through stochastic gradient ascend. Note that the log-posterior of the latent variable $Z$ in (4) can be fully decoupled as the summation of per-sample contribution as
@@
where $C_{2}$ is constant in $\theta$. To account for variation in model parameters instead of treating them as deterministic values, we can also employ a Bayesian neural network, which uses variational inference (VI) (Blei et al., 2017) to approximate (7). Specifically, a variational distribution $q_{\phi}(\theta)$ is introduced to approximate the true posteriors in (7) where $q_{\phi}(\theta)\sim\mathcal{N}(\theta|\mu_{\phi},\sigma_{\phi}^{2}I_{p})$. Note that $\phi=(\mu_{\phi},\sigma_{\phi}^{2})$ are learnable parameters for the variational distribution. The evidence lower bound (ELBO) for the posterior is given by
@@
where the first term represents the expected log-likelihood under the variational posterior, and the second term is the Kullback–Leibler divergence between the variational distribution and the prior over model parameters. To enable efficient gradient-based optimization with respect to the variational parameters ($\phi$), we adopt the reparameterization trick (Kingma et al., 2015), which is denoted as
@@
where $\epsilon\sim\mathcal{N}(0,I_{d})$. $\odot$ is the element-wise product. $d$ is the dimensionality of the parameter space of the generative model. Applying variational inference in BNNs using mini-batches can result in high-variance gradient estimates. To mitigate this issue, we adopt the Flipout technique (Wen et al., 2018) when implementating the reparameterization trick. Flipout reduces gradient variance by decorrelating model parameters perturbations across different training examples within the same mini-batch. Specifically, instead of using a single shared random perturbation for all examples, Flipout generates pseudo-independent perturbations for each data point in the mini-batch, thereby decorrelating the resulting gradients, reducing their variance, and improving training stability.
In each iteration, we first update the variational parameters by maximizing the ELBO in (8), and then draw samples of the model parameters according to (9). Conditioned on these sampled parameters, we can use a standard forward pass through the network, which includes linear transformations followed by nonlinear activation functions, to compute the mean and variance functions $\mu(Z)$ and $\Sigma(Z)$ or their derivatives. Importantly, each iteration requires only a randomly sampled mini-batch of observed data. Within each iteration, we first compute the gradient of (4) with respect to the latent variable $Z$ and update $Z$ for each individual using stochastic gradient ascent where an Adam (Kingma, 2014)) optimizer is used conditional on the current model parameters. We then compute the gradient of each ELBO term in (8) with respect to the variational parameters $\phi$ of the generative model, and update $\phi$ via stochastic gradient ascent to maximize the ELBO given the current latent variables.
@@ 2.4 Arbitrary Conditional Inference
we tackle the above integral (13) through two steps. In the first step, we draw multiple latent variable $Z$ from the posterior ditribution $p_{Z\mid X_{\mathcal{A}}}(Z\mid X_{\mathcal{A}})$ using Markov chain Monte Carlo (MCMC) reviewed in Geyer (2011). In the second step, we draw $X_{\mathcal{B}}$ from $p_{X_{\mathcal{B}}\mid Z,X_{\mathcal{A}}}(X_{\mathcal{B}}\mid Z,X_{\mathcal{A}})$ based on all Monte Carlo samples of $Z$ where $p_{X_{\mathcal{B}}\mid Z,X_{\mathcal{A}}}(X_{\mathcal{B}}\mid Z,X_{\mathcal{A}})$ has a closed form as a multivariate Gaussian distribution.
In the first step of sampling the posterior samples of latent variable $Z$, we adopt Hamiltonian Monte Carlo (HMC) algorithm, which uses gradient-informed dynamical trajectories to make long-distance proposals with high acceptance rates, achieving faster and better sampling efficiency (Neal and others, 2011). According to Bayes Theorem, the log posterior of latent variable $Z$ can be represented as
@@
where $C_{3}=-logP_{X_{\mathcal{A}}}(X_{\mathcal{A}})$ and $X_{\mathcal{A}}|Z\sim\mathcal{N}(\mu_{\mathcal{A}}(Z),\Sigma_{\mathcal{A}}(Z))$. Here, $\mu_{\mathcal{A}}(Z)$ and $\Sigma_{\mathcal{A}}(Z))$ are the corresponding partition components in the $\mu(Z)$ and $\Sigma(Z)$ for $X_{\mathcal{A}}$, respectively. Under the diagonal covariance simplification, $\Sigma_{\mathcal{A}}(Z)$ also has a diagonal structure. For each test data point for $X_{\mathcal{A}}$, we keep a sufficient number of MCMC posterior samples of latent variable $Z$. Note that sampling low-dimensional latent variable $Z$ for each test data point is fully decoupled, which enables efficient parallelization and further improving computational efficiency. We implemented MCMC sampling process fully on GPU in parallel by using TensorFlow Probability (TFP) library (Dillon et al., 2017) for acceleration.
@@
Through the above two-step sampling process, we can perform conditional inference by sampling from $P_{X_{\mathcal{B}}\mid X_{\mathcal{A}}}(X_{\mathcal{B}}|X_{\mathcal{A}})$. Let us denote the test data in inference stage as $\{X_{\mathcal{A}}^{i}|i=1,...,N_{test}\}$ where $N_{test}$ is the total test sample size. We obtain the posterior samples of latent variable through the first step of sampling process, which are denoted as $\{Z^{i,m}|i=1,...,N_{test},m=1,...,N_{MC}\}$ where $N_{MC}$ is the MCMC sample size. In the second step of sampling, all the posterior samples of latent variable $Z$ are fed to the generative model $G$ to obtain the mean and covariance functions in (15) in order to sample from (13). The final conditional posterior samples are denoted as $\{X_{\mathcal{B}}^{i,m}|i=1,...,N_{test},m=1,...,N_{MC}\}$. The point estimate for the prediction can be constructed based on the conditional mean of the posterior samples as

### v2

# An AI-powered Bayesian Generative Modeling Approach for Arbitrary Conditional Inference

@@ Abstract
Modern data analysis increasingly requires flexible conditional inference $P(X_{\mathcal{B}}\mid X_{\mathcal{A}})$, where $(X_{\mathcal{A}},X_{\mathcal{B}})$ is an arbitrary partition of the observed variables $X$. Existing approaches are either restricted to a fixed conditioning structure or depend strongly on the distribution of conditioning masks during training. To address these limitations, we introduce Bayesian generative modeling (BGM), a unified framework for arbitrary conditional inference. BGM learns a generative model of $X$ via a stochastic iterative Bayesian updating algorithm in which model parameters and latent variables are updated until convergence. Once trained, any conditional distribution can be obtained without retraining. Empirically, BGM achieves superior predictive performance with posterior predictive intervals, demonstrating that a single learned model can serve as a universal engine for conditional prediction with principled uncertainty quantification. We provide theoretical guarantees for convergence of the stochastic iterative algorithm, statistical consistency, and conditional risk bounds. The proposed BGM framework leverages modern AI to capture complex relationships among variables while adhering to Bayesian principles, offering a promising approach for a wide range of applications in modern data science. Code for BGM is available at https://github.com/liuq-lab/bayesgm. Document of BGM is available at https://bayesgm.readthedocs.io.
@@ 1 Introduction
Conditional inference concerns the problem of inferring the distribution of a response variable given a set of predictor variables, a fundamental task in both machine learning and statistical science (Reid 1995). Classical supervised learning typically focuses on predicting a response variable given a fixed set of predictors. However, many modern data analysis tasks require arbitrary conditional inference, where one seeks to estimate $P(X_{\mathcal{B}}\mid X_{\mathcal{A}})$ for any partition $(X_{\mathcal{A}},X_{\mathcal{B}})$ of the observed variables $X$. Such flexibility is essential in dynamic real-world scenarios with heterogeneous or time-varying observation pattern. Discriminative conditional prediction methods typically lack this flexibility as they cannot perform new conditional inference without modifying the model architecture or retraining from scratch when the conditioning set changes.
To enable arbitrary conditional inference, early classical approaches utilized linear latent variable models and finite mixtures, such as the Mixture of Factor Analyzers (MFA) (Ghahramani et al. 1996), which models $P(X)$ as a finite mixture of low-rank Gaussian components and thereby supports arbitrary conditional distributions in closed-form. To relax the assumption of a fixed number of mixture components, nonparametric Bayesian models such as Dirichlet process mixtures (DPMs) (Ferguson 1973, Antoniak 1974) place priors on infinite mixtures. While DPMs offer a more flexible framework for joint density estimation and arbitrary conditioning, the posterior computation remains challenging and scaling to large, high-dimensional datasets is nontrivial.
To better capture nonlinear relationships, non-linear latent variable models such as the Gaussian Process Latent Variable Model (GPLVM) (Lawrence and Hyvärinen 2005) and its Bayesian extensions (Bayesian GPLVM) (Titsias and Lawrence 2010) emerged as powerful alternatives. By placing a Gaussian Process prior on the mapping from a low-dimensional latent space to the observed data space, GPLVMs provide a flexible, nonparametric approach to modeling complex joint distributions. These methods face severe computational bottlenecks due to exact Gaussian Process inference, making it difficult to scale and apply to modern large-scale data where both the sample size and feature dimension are substantial.
Recent years have witnessed rapid progress in conditional inference driven by advances in generative artificial intelligence (AI). This has led to the development of AI models to handle varying conditionings, such as Arbitrary Conditional Evaluation (ACE) (Strauss and Oliva 2021), Variational Autoencoders with Arbitrary Conditioning (VAEAC) (Ivanov et al. 2019), and Arbitrary Conditional Flows (ACFlow) (Li et al. 2020). By training neural networks across random masking patterns or employing conditional normalizing flows, these models can capture highly nonlinear relationships and approximate arbitrary conditional prediction. However, these AI-driven approaches either heavily rely on the distribution of masking during training process or impose architectural constraints to neural networks. Furthermore, these methods focus primarily on accurate density estimation or point prediction. They often lack a coherent statistical mechanism for principled uncertainty quantification, which is critical in high-stakes domains where decisions must account for variability.
@@
To address this widespread lack of rigorous uncertainty quantification in predictive modeling, a separate line of research has focused on conformal prediction (CP), which provides a model-agnostic framework for constructing prediction sets with guaranteed finite-sample coverage (Vovk et al. 2005, Lei et al. 2018, Barber et al. 2021). Conformal prediction methods transform point predictors into valid predictive intervals or regions without relying on parametric assumptions, ensuring that the true response lies within the constructed set with a user-specified probability. Recent developments have substantially expanded the scope of conformal prediction methods, including quantile-based approaches (Romano et al. 2019), as well as localized conformal prediction (Guan 2023), which adapts coverage to heterogeneous data distributions. These CP approaches have become a cornerstone for uncertainty quantification in modern machine learning applications due to their simplicity and strong theoretical guarantees (Kato et al. 2023). However, while CP methods address the need for uncertainty quantification, they remain fundamentally constrained by the fixed conditioning structure. Furthermore, conformal prediction typically offers marginal coverage rather than full conditional calibration.
@@
To address above challenges, we propose a Bayesian Generative Modeling (BGM) framework that leverages the power of modern AI techniques while adhering to the foundational principles of Bayesian inference. BGM learns the underlying generative process of the observed variables from a low-dimensional latent space, estimated by a stochastic iterative updating algorithm. Once trained, BGM enables arbitrary conditional inference, predicting the distribution of any subset of variables given the remaining without retraining or modifying the model architecture. This “train once, infer anywhere” property allows BGM to naturally accommodate dynamically changing conditioning sets that commonly arise in practice. By adopting a coherent Bayesian principle, BGM delivers posterior predictive intervals at any user-specified significance level. The proposed BGM framework thus combines the flexibility of modern AI models with the statistical coherence of Bayesian inference, offering a powerful and scalable solution for conditional prediction and uncertainty quantification in high-dimensional settings. We summarize the contributions of BGM as follows.
  • BGM formulates arbitrary conditional inference as posterior updating in a AI-powered Bayesian latent variable model, providing a unified probabilistic framework that fundamentally extends existing conditional inference methods, which are typically limited to one fixed conditioning structure or require a training mask distribution.
  • We develop a stochastic iterative updating algorithm with convergence guarantee and establish posterior consistency and conditional risk bounds under mild regularity conditions. It also ensures efficient scaling to large, high-dimensional datasets since only a mini-batch of data is required at each step during training and posterior inference for each test sample is independent.
@@
  • We demonstrate that expressive neural parameterizations can be embedded within this Bayesian framework. Extensive experiments demonstrate that BGM achieves superior predictive accuracy with uncertainty quantification compared to leading conformal prediction methods and data imputation methods. These results highlight BGM’s practical value as a principled tool for conditional inference in modern data analysis.
@@ 2.1 Problem Setup
Although estimating the expected response $\mathbb{E}[X_{\mathcal{B}}|X_{\mathcal{A}}]$ remains central to most regression analysis, this statistic alone provides an incomplete picture of the underlying uncertainty. Practitioners often require much richer distributional information, such as conditional variances, quantiles, tail probabilities, or predictive intervals, to assess risk, reliability, and variability in predictions. Therefore, our focus extends beyond point estimation to the broader goal of accurately characterizing the full conditional distribution of interest, providing both point estimate and uncertainty quantification across arbitrary subsets of variables to condition on.
@@ 2.2 Generative Process of Observed Data
where $\theta$ is a common parameter to specify all the conditional distribution of $X$ given $Z$. The prior distributions of both $Z$ and $\theta$ are set to be multivariate normal distributions. By default, we model the conditional distribution of $X|Z$ as normal distributions for continuous variables and logistic regression for discrete variables. In a typical continuous setting, the generative processes are defined as follows
@@
where both mean and covariance matrix are learnable functions of latent variable $Z$ parameterized by $\theta$. In practice, we can further simplify the covariance matrix $\Sigma(Z)$ as a diagonal structure $\Sigma(Z)=diag(\sigma^{2}_{1}(Z),...,\sigma^{2}_{p}(Z))$. In this diagonal simplification, $X_{\mathcal{A}}$ is independent of $X_{\mathcal{B}}$ given the latent variable $Z$. A richer covariance structures can capture residual conditional dependence (see our Discussion). Note that the generative function $G$ provides both mean function $\mu(Z)$ and covariance function $\Sigma(Z)$, which is represented by a neural network with two output heads parametrized by $\theta$. To account for uncertainty or variation in model parameter space, we can also adopt Bayesian neural network (BNN), where model parameters are treated as random variables instead of deterministic values (Goan and Fookes 2020, Jospin et al. 2022).
@@ 2.3 Stochastic Iterative Updating Algorithm
where $C_{1}=C-\frac{p}{2}log(2\pi)$, $X_{i}$ is the $i$-th element in $X$, and $\mu_{i}(Z)$ is the $i$-th element in the mean vector $\mu(Z)$. We update latent variable $Z$ for each sample during the training stage by maximizing (4) through stochastic gradient ascent. Note that the log-posterior of the latent variable $Z$ in (4) can be fully decoupled as the summation of per-sample contribution as
@@
where $C_{2}$ is constant in $\theta$. To account for variation in model parameters instead of treating them as deterministic values, we can also employ a Bayesian neural network, which uses variational inference (VI) (Blei et al. 2017) to approximate (7). Specifically, a variational distribution $q_{\phi}(\theta)$ is introduced to approximate the true posteriors in (7) where $q_{\phi}(\theta)\sim\mathcal{N}(\theta|\mu_{\phi},\sigma_{\phi}^{2}I_{p})$. Note that $\phi=(\mu_{\phi},\sigma_{\phi}^{2})$ are learnable parameters for the variational distribution. The evidence lower bound (ELBO) for the posterior is given by
@@
where the first term represents the expected log-likelihood under the variational posterior, and the second term is the Kullback–Leibler divergence between the variational distribution and the prior over model parameters. To enable efficient gradient-based optimization with respect to the variational parameters ($\phi$), we adopt the reparameterization trick (Kingma et al. 2015), which is denoted as
@@
where $\epsilon\sim\mathcal{N}(0,I_{d})$. $\odot$ is the element-wise product. $d$ is the dimensionality of the parameter space of the generative model. Applying variational inference in BNNs using mini-batches can result in high-variance gradient estimates. To mitigate this issue, we adopt the Flipout technique (Wen et al. 2018) for implementation of the reparameterization trick. Flipout reduces gradient variance by decorrelating model parameters perturbations across different training examples within the same mini-batch. Specifically, instead of using a single shared random perturbation for all examples, Flipout generates pseudo-independent perturbations for each data point in the mini-batch, thereby decorrelating the resulting gradients, reducing their variance, and improving training stability.
In each iteration, we first update the variational parameters by maximizing the ELBO in (8), and then draw samples of the model parameters according to (9). Conditioned on these sampled parameters, we can use a standard forward pass through the network, which includes linear transformations followed by nonlinear activation functions, to compute the mean and variance functions $\mu(Z)$ and $\Sigma(Z)$ or their derivatives. Importantly, each iteration requires only a randomly sampled mini-batch of observed data. Within each iteration, we first compute the gradient of (4) with respect to the latent variable $Z$ and update $Z$ for each individual using stochastic gradient ascent where an Adam (Kingma 2014)) optimizer is used conditional on the current model parameters. We then compute the gradient of each ELBO term in (8) with respect to the variational parameters $\phi$ of the generative model, and update $\phi$ via stochastic gradient ascent to maximize the ELBO given the current latent variables.
@@ 2.4 Arbitrary Conditional Inference
we tackle the above integral (10) through two steps. In the first step, we draw multiple latent variable $Z$ from the posterior distribution $p_{Z\mid X_{\mathcal{A}}}(Z\mid X_{\mathcal{A}})$ using Markov chain Monte Carlo (MCMC) reviewed in Geyer (2011). In the second step, we draw $X_{\mathcal{B}}$ from $p_{X_{\mathcal{B}}\mid Z,X_{\mathcal{A}}}(X_{\mathcal{B}}\mid Z,X_{\mathcal{A}})$ based on all Monte Carlo samples of $Z$ where $p_{X_{\mathcal{B}}\mid Z,X_{\mathcal{A}}}(X_{\mathcal{B}}\mid Z,X_{\mathcal{A}})$ has a closed form as a multivariate Gaussian distribution.
In the first step of sampling the posterior samples of latent variable $Z$, we adopt Hamiltonian Monte Carlo (HMC) algorithm, which uses gradient-informed dynamical trajectories to make long-distance proposals with high acceptance rates, achieving faster and better sampling efficiency (Neal and others 2011). According to Bayes Theorem, the log posterior of latent variable $Z$ can be represented as
@@
where $C_{3}=-logP_{X_{\mathcal{A}}}(X_{\mathcal{A}})$ and $X_{\mathcal{A}}|Z\sim\mathcal{N}(\mu_{\mathcal{A}}(Z),\Sigma_{\mathcal{A}}(Z))$. Here, $\mu_{\mathcal{A}}(Z)$ and $\Sigma_{\mathcal{A}}(Z))$ are the corresponding partition components in the $\mu(Z)$ and $\Sigma(Z)$ for $X_{\mathcal{A}}$, respectively. Under the diagonal covariance simplification, $\Sigma_{\mathcal{A}}(Z)$ also has a diagonal structure. For each test data point for $X_{\mathcal{A}}$, we keep a sufficient number of MCMC posterior samples of latent variable $Z$. Note that sampling low-dimensional latent variable $Z$ for each test data point is fully decoupled, which enables efficient parallelization and further improving computational efficiency. We implemented MCMC sampling process fully on GPU in parallel by using TensorFlow Probability (TFP) library (Dillon et al. 2017) for acceleration.
@@
Through the above two-step sampling process, we can perform conditional inference by sampling from $P_{X_{\mathcal{B}}\mid X_{\mathcal{A}}}(X_{\mathcal{B}}|X_{\mathcal{A}})$. Let us denote the test data in inference stage as $\{X_{\mathcal{A}}^{i}|i=1,...,N_{test}\}$ where $N_{test}$ is the sample size in test set. We obtain the posterior samples of latent variable through the first step of sampling process, which are denoted as $\{Z^{i,m}|i=1,...,N_{test},m=1,...,N_{MC}\}$ where $N_{MC}$ is the MCMC sample size. In the second step of sampling, all the posterior samples of latent variable $Z$ are fed to the generative model $G$ to obtain the mean and covariance functions in (15) in order to sample from (13). The final conditional posterior samples are denoted as $\{X_{\mathcal{B}}^{i,m}|i=1,...,N_{test},m=1,...,N_{MC}\}$. The point estimate for the prediction can be constructed based on the conditional mean of the posterior samples as